May 20, 2022 Tech group SAD meeting

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place =

Friday May 20, 2022 10:30am Hall D Counting House and Teams


  • Safety - Electrical
  • Key validation due this month
  • Hall will most likely be locked up 2 weeks from today!
  • First wave of FCAL2 PR approved $17K for Support plates, $24K cooling blocks waiting for signature
  • Accelerator schedule - HCO in full gear
  • Hall D SAD Tasks
    • CPP INstall Completed
      • Adjust Lead absorber procedure
      • Change Target procedure
    • PS Parasitic Detectors stages - low priority
    • FCAL Insert Module fabrication
      • 60 brass straps per week needed now?
      • matching mu metal to steel shields
      • cutting tool
      • Moving tent end of June to EEL 108
    • Tagger NMR repair and HCO - Scot
    • Tagger controls/network switch UPS
    • Solenoid HCO - complete
    • Downstream vacuum pipe install - complete - turn pump back on in final HCO
      Lum mon.jpg
    • Add casters to Microscope stand
  • TEDF work area move in June to test lab high bay or ESB
  • FCAL Insert Drawings - 98% complete
    • need cooling pipe to chiller assy
    • need dark room mod for electronic patch panels
    • need to complete 0200 final assy dwg
    • need installation drawing
    • Procurement plan developed
  • New Profiler install - not until August - new plates arrived
  • New 3x3 CCAL for TAC - cabling to be done
    TAC-3x3 in.jpg
    3x3 in.jpg
  • DIRC water skid and fiber hookup
  • Controls
    • FDC chiller cooling water flow -
    • FDC start up procedure with set point error - restarted yesterday, any issues?
  • HCO started...
  • General Maintenance
  • On Call - Keith
  • ePAS Training and trial
  • Tidy Friday? Thorough cleaning next week?

Action Items

  • Complete HCO
  • Get Keith fully trained on Brass straps
  • DIRC water skid start up